# BlockLLM: Memory-Efficient Adaptation of LLMs by Selecting and Optimizing the Right Coordinate Blocks

### v1

@@ 3 Experiments
We evaluated BlockLLM on both finetuning tasks and pre-training tasks ***All our experiments were based on the code released by [34]. We thank the authors for releasing their code and clearly documenting their experiment setup in the paper.. The finetuning experiments were conducted in Tesla V100 gpu ($32$ GB) and the pretraining tasks were conducted on NVIDIA A40 ($48$ GB) and A100 GPUs ($80$ GB) (one GPU per experiment).
3.1 Finetuning on GLUE
The General Language Understanding Evaluation (GLUE) benchmark [31] is widely used to evaluate the performance of NLP models across a range of tasks, including sentiment analysis, question answering, and textual entailment. We benchmarked the performance of BlockLLM against GaLore [34] and full finetuning (FFT) using the pre-trained RoBERTa model [17] on GLUE tasks. We did not compare against LoRA [9] and its variants [16, 11], because our experimental setup aligns with those described in [16, 34] and the results for these methods are already documented in [34]. For BlockLLM, we conducted hyperparameter tuning for the learning rate, and our experimental setup is detailed in A.5. For GaLore, we used the learning rate specified in [34]. In our experiments, we used $s=0.95$. We evaluated both performance and memory consumption during the training process for all methods. The results, presented in Tables 2 and 3, indicate that BlockLLM outperforms the other models in all tasks while achieving approximately a 13.5% reduction in memory usage on average.
MRPC COLA STS-B RTE SST2 MNLI QNLI QQP
Block-LLM $\mathbf{3.97}$ $\mathbf{2.8}$ $\mathbf{3.48}$ $\mathbf{9.1}$ $\mathbf{3.6}$ $\mathbf{13.7}$ $\mathbf{12.8}$ $8.36$
GaLore (rank=8) $4.52$ $4.2$ $4.8$ $9.7$ $3.87$ $15.1$ $14.8$ $8.43$
GaLore (rank=4) $4.52$ $4.2$ $4.8$ $9.7$ $3.86$ $15.2$ $14.8$ $\mathbf{8.03}$
FFT $4.24$ $3.67$ $4.4$ $10.28$ $3.82$ $15.53$ $15.02$ $9.22$
MRPC COLA STS-B RTE SST2 MNLI QNLI QQP
Block-LLM 91.8 $\mathbf{63.8}$ 90.02 80.14 $\mathbf{94.95}$ $\mathbf{87.75}$ 92.95 91.36
GaLore (rank=8) 89.96 62.5 $\mathbf{91.1}$ 79.78 94.38 87.17 $\mathbf{92.97}$ 91.11
GaLore (rank=4) 91.7 61.67 91.09 79.78 94.04 87 92.65 91.06
FFT $\mathbf{92.36}$ 62.84 91.1 $\mathbf{80.5}$ 94.57 87.18 92.33 $\mathbf{92.28}$
Table 3: Score Comparison Across Different GLUE Tasks
3.2 Pre-training on Llama model
@@
We also compared BlockLLM with GaLore [34] in pre-training LLaMA-based large language models [30] on the C4 (Colossal Clean Crawled Corpus) dataset [23]. The C4 dataset is a large-scale, cleaned version of the Common Crawl web corpus used for pre-training language models, featuring diverse and high-quality text from the internet. Our experiment setup is similar to [34], following the setup from [16]. We fine-tuned the learning rate for BlockLLM while keeping all other hyperparameters fixed across experiments. We ran BlockLLM with the experimental setup described in A.6 and computed the perplexity scores from final evaluation loss and maximum memory usage in GB(Gigabytes). We ran GaLore for $10\%$ of total iterations to observe memory consumption, and used the results from [34] for comparison. The perplexity and the memory are shown in the Figure 4 and Table 4.
@@
Table 4: Comparison of Accuracy and VRAM Memory (GB) of LLama Models with GaLore. Note that for GaLore, we used AdamW as the underlying optimizer.
We note here that BlockLLM takes a little bit longer to converge for some experiments ($130m$ and $350m$ models) compared to full parameter learning methods such as GaLore in the pre-training experiment. We suspect this is due to the fact that we are dealing with noisy gradients in the earlier iterations of training. However, BlockLLM converges to the state-of-the-art perplexity score in a few more iterations compared to GaLore. This is issue is not present in the finetuning experiments.
Effect of sparsity $s$. Here, we compare BlockLLM with sparsity values $s=0.5,0.7$ and $0.9$ against full parameter training $(s=0)$ and GaLore [34]. The results are presented in Figure 4. We observe that with $s=0.5$, BlockLLM consumes about $1.5$ GB less memory than Galore and higher sparsity values further reduce memory usage though this comes with the trade-off of requiring more training iterations for similar performance.
@@ 3.3 Ablation on Parameter Selection Strategy
We investigate whether selecting parameters with higher gradient norms, which are less frequently chosen, benefits training. We observed the training performance of the LLaMA 60M model on the C4 dataset while employing the strategy of randomly selecting block parameters for updates. Specifically, let $L=[l_{1},l_{2},\dots,l_{n}]$ represent the list of all layers. At iteration $t$, we randomly select $k$ layers from $L$ with equal probability until the required number of parameters is met according to the sparsity $s$. Let $D\subseteq L_{\text{shuffled}}$ be the subset of selected layers, defined as
$D=\{l_{i}\in L_{\text{shuffled}}:\sum_{l_{i}\in D}\text{param\_count}(l_{i})\leq s\times\sum_{l_{i}\in L}\text{param\_count}(l_{i})\}$
Figure 5: Comparison of random vs greedy BlockLLM on Llama 60m model with 50% parameters chosen.
[…]
Our observations indicate that BlockLLM converges faster than the random parameter selection method. Intuitively, this suggests that parameters with higher gradient norms may facilitate more rapid learning. One possible explanation is that parameters with higher gradient norms contribute to larger updates, thereby playing a more significant role in the training process during each iteration. Consequently, these parameters might accelerate the model’s convergence by making more substantial contributions to the optimization trajectory.
@@
To assess this, we conducted an experiment where parameters were selected based solely on their gradient norms. Specifically, for each layer $l\in L$, we computed the processed gradients $\tilde{G}_{l}$. We then sorted the layers in descending order based on the their processed gradient norms, $||\tilde{G}_{l}||$. From this ordered list, we selected the top $k$ layers that satisfied the sparsity requirement $s$. The results of this experiment are presented in Figure 7. We observed that when parameters were selected solely based on $||\tilde{G}||$, without considering the parameter visit frequency $f_{l}$ for any layer $l\in L$, the loss converged more slowly and to a higher value compared to BlockLLM.
@@ 4 Conclusion & Future work
In this paper, we introduced BlockLLM, a novel method for efficiently training large language models. By dynamically estimating and updating the importance of parameters during training, BlockLLM effectively achieves state-of-the-art performance while significantly reducing the memory footprint. Our method achieves the highest validation accuracy on GLUE finetuning tasks, sometimes even surpassing full finetuning. One key aspect of BlockLLM is that it does not presuppose the importance of layers but continuously evaluates and updates parameter importance throughout training. This adaptive approach allows for more flexible and efficient optimization compared to methods that assume certain parameters are critical from the outset. Additionally, BlockLLM preserves the original architecture without altering the model structure or restricting the parameter search space, making it suitable for various LLMs and tasks.
Broader Impacts
Our work aims to reduce the memory and computational requirements of training LLMs. First, our technology democratizes access to LLM training, making it more feasible for student researchers and institutions with limited computational resources to participate in cutting-edge AI research. Furthermore, the low-memory requirements of our method means that one can train with larger batch sizes and achieve faster convergence. This has a direct effect on the environment.
Future works.
Future work on BlockLLM could explore several promising avenues. Currently, our research has focused on parameter selection based on gradient norms, but BlockLLM can be seen as a framework for parameter-efficient training rather than a single algorithm. This opens the door to investigating alternative criteria for parameter selection, potentially tailored to specific problems or tasks. Moreover, while our ablation studies on BlockLLM’s hyperparameters have provided insights into their impact on training, further research is needed to understand how different layers might be affected by greedy parameter selection strategies. BlockLLM also complements existing memory-optimized training techniques, including those discussed in this paper. Exploring the integration of BlockLLM with methods like quantization or GaLore could further reduce memory consumption.
Appendix A Appendix / supplemental material
@@ A.1 Sparsity Accuracy Tradeoff
We performed magnitude pruning on the IMDb pre-trained model [19] weights at various sparsity levels and fine-tuned these pruned models on the GLUE-CoLA dataset [31]. The results of these experiments, detailing the relationship between sparsity and accuracy, are summarized in Table 5.
@@ A.2 Analysis of weight magnitudes
In this experiment, we pretrain DistilBERT on the IMDB dataset [19] and then finetune it on GLUE-CoLA [31] with sparsity $s=0.7$. We then plot the histogram of the weight magnitudes $W^{t}$ where $\delta=|w_{i}^{0}-w_{i}^{t}|>\eta$, with $\eta$ as the threshold. We set $\eta=0.001$ in this case.
@@ A.3 Analysis of reduced parameter training
Here, we fine-tuned the GLUE datasets [31] on the DistilBERT model [27] pretrained on the IMDb dataset [19]. We varied the sparsity $s$ and update frequency $m$ while monitoring the number of unique parameters updated $q$. We ran the GLUE-SST2 experiments for $8400$ iterations and GLUE-STSB for $20000$ iterations. Additionally, we tracked the VRAM usage for the GLUE-SST2 dataset to compare it with the memory consumption of full-parameter fine-tuning, which is $7.9$ GB. This comparison aims to determine if reduced parameter training effectively decreases memory usage.
@@ A.4 VRAM memory
All memory values presented in our tables represent actual observed memory usage in gigabytes (GB) rather than estimates. Memory consumption was monitored using the "nvidia-smi" command, and the maximum memory usage recorded during the training process was noted.
@@ A.5 Finetuning on GLUE
In this section, we detail the experimental setup employed for our study. We fine-tuned the pre-trained RoBERTa model [17] , available in the HuggingFace library, on the GLUE benchmark tasks [31]The batch size was set to $32$ for the CoLA dataset, and $16$ for all other datasets. For all tasks, we used $s=0.95$ and $m=\frac{1}{4}\times$ total number of iterations. VRAM memory usage was monitored and recorded as described in Section A.4. The learning rates for the different tasks are as follows.
@@
A.6 Pre-training on Llama
We present the hyperparameters utilized for training the LLama models with sizes $60$M, $130$M and $350$M in 9. For $60$M and $130$M experiments, the maximum sequence length was set to $256$ with a gradient accumulation of $2$,and for $350$M with a batch size of $128$ with a gradient accumulation of $4$. A cosine annealing schedule was employed for learning rate adjustment, decaying to $10$% of the initial learning rate. For BlockLLM, no learning rate warmup was applied. However, for GaLore, the learning rate was warmed up for the first $10\%$ of training, following the approach outlined in [34]. The parameter $m$ was set to $50$ for all the experiments.

### v2

@@ 3 Experiments
We evaluated BlockLLM on both finetuning and pretraining tasks. [^1] All our experiments were based on the code released by Zhao et al. (2024). We thank the authors for making their code publicly available and for clearly documenting their experimental setup. The large-scale finetuning experiments were conducted on an H100 GPU (80 GB), while the pretraining tasks were performed on NVIDIA A40 (48 GB) and A100 GPUs (80 GB) (one GPU allocated per experiment). The remaining experiments were conducted using a Tesla V100 GPU (32 GB).
3.1 Large scale finetuning
We conducted a series of experiments to evaluate the effectiveness of our approach in finetuning large language models. Specifically, we utilized the LLaMA-2 model (Touvron et al., 2023) with 7 billion parameters, and fine-tuned it on the Alpaca dataset (Peng et al., 2023). Alpaca dataset Peng et al. (2023) is a widely used benchmark for instruction following tasks. It consists of diverse instruction-response pairs across multiple domains. For the finetuning process, we employed the Llama-factory framework (Zheng et al., 2024). Llama-factory (Zheng et al., 2024) is an open-source framework designed for efficient fine-tuning, inference, and deployment of LLaMA models.
We finetuned the LLaMA-2 model (Touvron et al., 2023) with the experimental setup as detailed in the Appendix 9. Gradient checkpointing was enabled in all cases. We compared the performance of BlockLLM against other state-of-the-art memory efficient training methods such as GaLore (Zhao et al., 2024), LoRA (Hu et al., 2021) and BAdam (Luo et al., 2024). We excluded Adam (Kingma & Ba, 2014) from the comparison as it exceeds the available memory on an 80GB GPU, with an estimated requirement of over 120GB. We evaluated all methods based on peak memory consumed during training in GB(Gigabytes), training time, and both training and evaluation loss. The results are provided in the Figure 5.
Figure 5: The figures, arranged from left to right and top to bottom, compare training loss, evaluation loss, peak memory usage, and training time for BlockLLM, LoRA, BAdam, and GaLore. BlockLLM demonstrates superior performance, with lower memory usage and reduced training time.
The results indicate that BlockLLM converges to a lower loss value than the other methods, while requiring substantially less memory. Additionally, BlockLLM demonstrates strong generalization, achieving the lowest evaluation loss across all methods. In terms of runtime, BlockLLM performs comparably to BAdam and is faster than the other two baselines.
3.2 Pretraining on Llama model
@@
We also compared BlockLLM with GaLore (Zhao et al., 2024) in pretraining LLaMA-based large language models Touvron et al. (2023) on the C4 (Colossal Clean Crawled Corpus) dataset (Raffel et al., 2020). The C4 dataset is a large-scale, cleaned version of the Common Crawl web corpus used for pre-training language models, featuring diverse and high-quality text from the internet. Our experiment setup is similar to Zhao et al. (2024), following the setup from Lialin et al. (2023). We finetuned the learning rate for BlockLLM while keeping all other hyperparameters fixed across experiments. We ran BlockLLM with the experimental setup described in A.7 and computed the perplexity scores from final evaluation loss and maximum memory usage in GB(Gigabytes). We ran GaLore for $10\%$ of total iterations to observe memory consumption, and used the results from Zhao et al. (2024) for comparison. The perplexity and the memory are shown in Figure 6 and Table 1.
@@
Table 1: Comparison of BlockLLM’s accuracy and VRAM memory usage (GB) for LLaMA models with GaLore. BlockLLM demonstrates reduced memory consumption while maintaining comparable performance.
We note here that BlockLLM takes a little bit longer to converge in some pretraining experiments ($130m$ and $350m$ models) compared to GaLore in the pretraining experiment. We suspect this is due to the fact that we are dealing with noisy gradients in the earlier iterations of training. However, BlockLLM converges to the state-of-the-art perplexity score in a few more iterations compared to GaLore. This issue is not present in the finetuning experiments.
Effect of sparsity $s$. Here, we compare BlockLLM with sparsity values $s=0.5,0.7$ and $0.9$ against GaLore (Zhao et al., 2024). The results are presented in Figure 6. We observe that with $s=0.5$, BlockLLM consumes about $1.5$ GB less memory than Galore and higher sparsity values further reduce memory usage though this comes with the trade-off of requiring more training iterations for similar performance.
@@ 3.3 Ablation on Parameter Selection Strategy
The parameter selection strategy is the cornerstone of our method, where we hypothesize that layers with large $||\tilde{G}_{l}||$ are important for training. To validate this, we conducted an ablation study in which we deliberately chose parameters with small $||\tilde{G}_{l}||$ (opposite to BlockLLM). Specifically, we finetuned LLaMA2 7B model on the Alpaca dataset, where parameters with the smallest gradient norms were selected for updates. We call this method BLockLLM-SubOPT. Naturally, we expect that this parameter selection strategy to be severely detrimental to the training process.
Let $L=\{l_{1},l_{2},\dots,l_{n}\}$ represent the set of all layers. Now, for each layer $l\in L$, we computed the processed gradients $\tilde{G}_{l}$. We then sorted the layers in ascending order based on the their processed gradient norms $||\tilde{G}_{l}||$ and computed $||\tilde{G}_{l}||/f_{l}$. From this ordered list, we selected top $k$ layers with small gradient norms until the sparsity requirement $s$ is satisfied. Then, in each iteration $t$, the optimizer updates only the parameters in the selected $k$ layers. We conducted hyperparameter tuning and set $m=100$ for $s=0.95$. Now, we compared the training loss of both the methods and the comparative results are presented in Figure 7. As expected, the results show that BLockLLM-SubOPT exhibits significantly higher training loss and converges slower than BlockLLM.
@@
To evaluate this, we conducted a similar experiment on the LLaMA 60M model pretrained on the C4 dataset, where we assessed BlockLLM’s strategy to select layers based on layer visit frequency $f$. To perform this, we selected layers solely based on their gradient norms without considering $f$. We compared this method with BlockLLM and the results are presented in Figure 7. Our hypothesis was that selecting parameters solely based on $||\tilde{G}||$ might result in higher loss during the initial iterations, followed by a gradual reduction in later iterations. This could be because of the noisy gradients early in training. Important layers might not be chosen resulting in higher loss compared BlockLLM. As expected, we observed similar behavior as shown in the figure 7.
Figure 7: Ablation study on parameter selection criteria. The figure on the left illustrates the advantage of selecting parameters based on gradient norms and the figure on the right demonstrates the benefit of incorporating layer selection frequency.
@@ 4 Conclusion & Future work
In this paper we introduced BlockLLM, a novel method for efficiently training large language models. By dynamically estimating and updating the importance of parameters during training, BlockLLM effectively achieves state-of-the-art performance while significantly reducing the memory footprint. Our method achieves the highest validation accuracy on the finetuning tasks, sometimes even surpassing full finetuning. One key aspect of BlockLLM is that it does not presuppose the importance of layers but continuously evaluates and updates parameter importance throughout training. This adaptive approach allows for more flexible and efficient optimization compared to methods that assume certain parameters are critical from the outset. Additionally, BlockLLM preserves the original architecture without altering the model structure or restricting the parameter search space, making it suitable for various LLMs and tasks.
Broader Impacts Our work aims to reduce the memory and computational requirements of training LLMs. First, our technology democratizes access to LLM training, making it more feasible for student researchers and institutions with limited computational resources to participate in cutting-edge AI research. Furthermore, the low-memory requirements of our method means that one can train with larger batch sizes and achieve faster convergence. This has a direct effect on the environment.
Future works. Future work on BlockLLM could explore several promising avenues. Currently, our research has focused on parameter selection based on gradient norms, but BlockLLM can be seen as a framework for parameter-efficient training rather than a single algorithm. This opens the door to investigating alternative criteria for parameter selection, potentially tailored to specific problems or tasks. Moreover, while our ablation studies on BlockLLM’s hyperparameters have provided insights into their impact on training, further research is needed to understand how different layers might be affected by greedy parameter selection strategies. BlockLLM also complements existing memory-optimized training techniques, including those discussed in this paper. Exploring the integration of BlockLLM with methods like quantization or GaLore could further reduce memory consumption.
Appendix A Appendix
@@ A.1 Sparsity Accuracy Tradeoff
We performed magnitude pruning on the IMDb pre-trained model (Maas et al., 2011) weights at various sparsity levels and fine-tuned these pruned models on the GLUE-CoLA dataset (Wang et al., 2018). The results of these experiments, detailing the relationship between sparsity and accuracy, are summarized in Table 2.
@@ A.2 Analysis of weight magnitudes
In this experiment, we pretrain DistilBERT on the IMDB dataset (Maas et al., 2011) and then finetune it on GLUE-CoLA (Wang et al., 2018) with sparsity $s=0.7$. We then plot the histogram of the weight magnitudes $W^{t}$ where $\delta=|w_{i}^{0}-w_{i}^{t}|>\eta$, with $\eta$ as the threshold. We set $\eta=0.001$ in this case.
@@ A.3 Analysis of reduced parameter training
[…]
We conducted this experiment using the same setup: finetuning the DistilBERT model on the GLUE-CoLA dataset after pretraining on the IMDb dataset. The results of this experiment are summarized in the Table 3.
$1-s$ $q$ $m$ Accuracy Matthews Correlation
$0.1$ $0.58$ $1000$ $82.55$ $0.5882$
$0.02$ $0.36$ $1000$ $82.45$ $0.5711$
$0.02$ $0.14$ $4000$ $82.45$ $0.5794$
$0.02$ $0.10$ $6000$ $82.07$ $0.5679$
Table 3: Impact of Update Frequency and Sparsity on CoLA Dataset
Next, we fine-tuned the GLUE datasets (Wang et al., 2018) on the DistilBERT model (Sanh et al., 2019) pretrained on the IMDb dataset (Maas et al., 2011). We varied the sparsity $s$ and update frequency $m$ while monitoring the number of unique parameters updated $q$. We ran the GLUE-SST2 experiments for $8400$ iterations and GLUE-STSB for $20000$ iterations. Additionally, we tracked the VRAM usage for the GLUE-SST2 dataset to compare it with the memory consumption of full-parameter fine-tuning, which is $7.9$ GB. This comparison aims to determine if reduced parameter training effectively decreases memory usage.
@@ A.4 VRAM memory
All memory values presented in our tables represent actual observed memory usage in gigabytes (GB) rather than estimates. Memory consumption was monitored using the “nvidia-smi” command, and the maximum memory usage recorded during the training process was noted.
@@ A.5 Finetuning on GLUE
The General Language Understanding Evaluation (GLUE) benchmark (Wang et al., 2018) is widely used to evaluate the performance of NLP models across a range of tasks, including sentiment analysis, question answering, and textual entailment. We benchmarked the performance of BlockLLM against GaLore (Zhao et al., 2024) and full finetuning (FFT) using the pre-trained RoBERTa model (Liu et al., 2019) on GLUE tasks. We did not compare against LoRA Hu et al. (2021) and its variants Lialin et al. (2023); Kamalakara et al. (2022), because our experimental setup aligns with those described in ReLoRA and GaLore(Lialin et al., 2023; Zhao et al., 2024). The results for these methods are already documented in GaLore(Zhao et al., 2024).
For BlockLLM, we conducted hyperparameter tuning for the learning rate and the learning rates for the different tasks are as follows.
@@
The batch size was set to $32$ for the CoLA dataset, and $16$ for all other datasets. For all tasks, we used $s=0.95$ and $m=\frac{1}{4}\times$ total number of iterations. VRAM memory usage was monitored and recorded as described in Section A.4. For GaLore, we used the learning rate specified in their original work (Zhao et al., 2024). In our experiments, we used $s=0.95$. We evaluated both performance and memory consumption during the training process for all methods. The results, presented in Tables 7 and 8, indicate that BlockLLM outperforms the other models in all tasks while achieving approximately a 13.5% reduction in memory usage on average.
MRPC COLA STS-B RTE SST2 MNLI QNLI QQP Avg.
Block-LLM $\mathbf{3.97}$ $\mathbf{2.8}$ $\mathbf{3.48}$ $\mathbf{9.1}$ $\mathbf{3.6}$ $\mathbf{13.7}$ $\mathbf{12.8}$ $8.36$ $\mathbf{7.2}$
GaLore (rank=8) $4.52$ $4.2$ $4.8$ $9.7$ $3.87$ $15.1$ $14.8$ $8.43$ $8.18$
GaLore (rank=4) $4.52$ $4.2$ $4.8$ $9.7$ $3.86$ $15.2$ $14.8$ $\mathbf{8.03}$ $8.14$
FFT $4.24$ $3.67$ $4.4$ $10.28$ $3.82$ $15.53$ $15.02$ $9.22$ $8.27$
MRPC COLA STS-B RTE SST2 MNLI QNLI QQP Avg.
Block-LLM 91.8 $\mathbf{63.8}$ 90.02 80.14 $\mathbf{94.95}$ $\mathbf{87.75}$ 92.95 91.36 $\mathbf{86.6}$
GaLore (rank=8) 89.96 62.5 $\mathbf{91.1}$ 79.78 94.38 87.17 $\mathbf{92.97}$ 91.11 $86.12$
GaLore (rank=4) 91.7 61.67 91.09 79.78 94.04 87 92.65 91.06 $86.12$
FFT $\mathbf{92.36}$ 62.84 91.1 $\mathbf{80.5}$ 94.57 87.18 92.33 $\mathbf{92.28}$ $\mathbf{86.6}$
Table 8: Score Comparison Across Different GLUE Tasks
A.6 Large scale finetuning of Llama 2 on alpaca
We provide more details on the hyperparameters used in training the Llama 2 model Touvron et al. (2023) on Alpaca datasetPeng et al. (2023).
Hyperparameter BlockLLM GaLore LoRA BAdam
Learning Rate (LR) $1\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$
LR Scheduler Cosine (lr_min = 0) Cosine (lr_min = 0) Cosine (lr_min = 0) Cosine (lr_min = 0)
Epochs 3 3 3 3
Batch Size 8 8 8 8
Gradient Accumulation 2 2 2 2
Weight Decay 0.01 0.01 0.01 0.01
Sparsity ($s$) 0.95 NA NA NA
Patience(m) 100 NA NA NA
K NA NA NA 100
rank NA 8 8 NA
$\alpha$ NA 2 $4\times rank$ NA
Table 9: Hyperparameter Details for finetuning LLaMA 2 Alpaca dataset
A.7 Pretraining on Llama
We present the hyperparameters utilized for training the LLama models with sizes $60$M, $130$M and $350$M in 10. For $60$M and $130$M experiments, the maximum sequence length was set to $256$ with a gradient accumulation of $2$,and for $350$M with a batch size of $128$ with a gradient accumulation of $4$. A cosine annealing schedule was employed for learning rate adjustment, decaying to $10$% of the initial learning rate. For BlockLLM, no learning rate warmup was applied. However, for GaLore, the learning rate was warmed up for the first $10\%$ of training, following the approach outlined in Zhao et al. (2024). The parameter $m$ was set to $50$ for all the experiments.